Sofia is 5 times as old as her cousin. Four years from now she will be 3 times as old as her cousin. Whats the ages of both Sofia and her cousin?
Answer by Deina(147)

Variables Used

  • S = Sarah

  • c = cousin

What do we know? We know that:
S = 5c
S + 4 = 3(c + 4)

Substitute "5c" for S in the second equation:
5c + 4 = 3(c + 4)

Distribute:
5c + 4 = 3c + 12

Subtract 3c from each side:
5c -3c + 4 = 12

Subtract 4 from each side:
2c = 8

Divide both sides by 2 and find that
c = 4

Substitute "4" for "c" in the 1st equation:
S = 5(4) and we find that
Sarah is 20 years old and her cousin is 4 years old
& bob's your uncle!